OpenStudy (anonymous):

1/2(x + 4) + 3 = 2x(9/2 + 3/4) what is the value of x?

OpenStudy (pfenn1):

\[\frac 1 2 (x+4)+3=2x \left( \frac 9 2 + \frac 3 4 \right)\] rearrange and solve for x